Jerry Rice and Grain Stores has $4,880,000 in yearly sales. The firm earns 3 percent on each dollar of sales and turns over its assets 2.8 times per year. It has $196,000 in current liabilities and $359,000 in long-term liabilities. a. What is its return on stockholders equity? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) Return on stockholders equity b. If the asset base remains the same as computed in part a, but total asset turnover goes up to 3.40, what will be the new return on stockholders equity? Assume that the profit margin stays the same as do current and long-term liabilities. (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) New return on stockholders equity
Answer #1

Answer

  • All working forms part of the answer
  • Workings

A

Sales

$4,880,000

B

Net Income % on sales

3%

C = A x B

Net Income

$146,400

A

Sales

$4,880,000

B

Asset Turnover ratio

2.8

C = A/D

Total Assets

$1,742,857

D

Current Liabilities

$196,000

E

Long Term Liabilities

$359,000

F = C - D - E

Total Stockholder's Equity

$1,187,857

  • Requirement ‘a’

A

Net Income

$146,400

B

Total Stockholder's Equity

$1,187,857

C = A/B

Return on Stockholder's Equity

12.32% = Answer

  • Requirement ‘b’

A

Total Assets

$1,742,857

B

Asset Turnover ratio

3.4

C = A x B

New total Sales

$5,925,714

D

Net Income % on sale

3%

E = C x D

Net Income

$177,771

F [already calculated above, stays same]

Total Stockholder's Equity

$1,187,857

G = E/F

New Return on Stockholder's Equity

14.97% = Answer